OpenStudy (anonymous):

-8 Square root 5t to the 5th, multiplied by 5 Square root 3t

OpenStudy (campbell_st):

so it looks like \[-8\sqrt{5t^5} \times 5\sqrt{3t}\] which becomes \[-8 \times 5 \times \sqrt{5 \times 3 \times t^5 \times t} = -40\sqrt{15t^6}\] which can me written as\[-40 \sqrt{15(t^3)^2}\] you just need to finish the simplification

OpenStudy (anonymous):

\[\left( -8\sqrt{40t^{5}} \right)\left( 5\sqrt{15t} \right)\] Please help!!! I cannot figure this thing out!